OpenStudy (jjuden):

Solve using the perfect squares method: (2x+5)2=25

mathslover (mathslover):

Take square root both sides... \(\sqrt{(2x+5)^2} = \pm \sqrt{25}\)

mathslover (mathslover):

Can you simplify it ?

OpenStudy (jadeishere):

There's a sipler way to do this @mathslover ...

OpenStudy (jadeishere):

*simpler

mathslover (mathslover):

@Jadeishere - Yep sure! Please put up your method here. I'm glad that you put it up like this.

OpenStudy (jadeishere):

Well, in a standard form quadratic equation, you can simplify it to get (2x+5)2=25 but then you could simplify THAT to get (2x + 5)^2 = 5^2 the easiest way to go about this is 2x + 5 = 25 or 2x + 5 = -5

OpenStudy (jadeishere):

My teacher showed me that when I was in 8th grade.

OpenStudy (jjuden):

so how would i finish cause my teacher never taught me that

OpenStudy (jadeishere):

well then just isolate the variable, like you would any other equation. By the way, I meant 2x + 5 = 5 not 25

OpenStudy (jjuden):

ok

mathslover (mathslover):

That is what I going to do too @Jadeishere - \(\sqrt{(2x+5)^2} = \pm \sqrt{25}\) So, 2x + 5 = 5 or -5. The one showed by your teacher is not having the step of square rooting both sides which is okay for the beginners.

OpenStudy (jadeishere):

Oh, sorry >.<